A 200-turn solenoid having a length of 25 cm and diameter of 10 cm carries a current of 0.30 A. What is the magnitude of the magnetic field B near the center of the solenoid?


1
Expert's answer
2021-11-04T10:18:15-0400

The magnetic fields is given as follows:


"B = \mu_0\dfrac{NI}{L}"

where "\mu_0 = 4\pi\times 10^{-7}H/m" is the magnetic constant, "N = 200" number of turns, "I = 0.30A" is the current, "L = 25cm = 0.25m" is the length of the solenoid. Thus, obtain:


"B = 4\pi\times 10^{-7}\dfrac{H}{m}\cdot \dfrac{200\cdot 0.30A}{0.25m} \approx 3.0\times 10^{-4}T"

Answer. "3.0\times 10^{-4}T".